OpenStudy (anonymous):

What is the 6th term of the geometric sequence where a1 = −625 and a2 = 125? @IMStuck

OpenStudy (anonymous):

−0.2 0.2 −0.04 0.04

OpenStudy (jdoe0001):

so.... what's the "common ratio" you think? that is, the amount we multiplied the 1st term, to get the 2nd one?

OpenStudy (anonymous):

Uh. ;-;

OpenStudy (jdoe0001):

hmmm do you know what a geometric sequence is?

OpenStudy (anonymous):

Yes.

OpenStudy (jdoe0001):

so.... there's a multiplier to get to the 2nd term... so \(\large \begin{array}{ccllll} -625& 125&...\\ &\uparrow \\ &-625\cdot {\color{brown}{ r}} \end{array}\) so... what's that number "r' or ratio.... in order to get 125?

OpenStudy (anonymous):

Uh. 1/5?

OpenStudy (jdoe0001):

anyhow... to make it short.... the simpler way is just to keep in mind that \(\bf -625\cdot {\color{brown}{ r}}=125\implies {\color{brown}{ r}}=\cfrac{125}{-625}\)

OpenStudy (anonymous):

For the Whole answer i got 0.2 :c Can yew check if im right?

OpenStudy (anonymous):

I Kinda lost my calculator o-o

OpenStudy (jdoe0001):

well \(\bf a_{\color{blue}{ 6}}=a_1\cdot {\color{brown}{ r}}^{{\color{blue}{ 6}}-1}\to -625\cdot {\color{brown}{ \left(\frac{1}{5}\right)}}^{5}\to -\cfrac{1}{5}\to -0.2\) you're correct

OpenStudy (jdoe0001):

https://desmos.com/calculator <--- btw

OpenStudy (anonymous):

OH YAAAY. Can you help with a few more?

OpenStudy (jdoe0001):

you can always post anew... so if I dunno.. someone else may and we can revise each other